# Novel inner shadows of the Kerr black hole illuminated by a tilted thin accretion disk

### v2

@@ Abstract
The inner shadow of a black hole, as a projection of the event horizon, is regarded as a potential tool for testing gravitational theories and constraining system parameters. Whether this holds in the case of a tilted accretion disk warrants further investigation. In this paper, we employ a ray-tracing algorithm to simulate images of the Kerr black hole illuminated by a tilted thin accretion disk, with particular attention to the relationship between the inner shadow and system parameters. Our findings reveal that in the case of an equatorial accretion disk, the Kerr black hole exhibits a minimum inner shadow size of $S_{\textrm{min}}=13.075$ M2, where M denotes the black hole mass. This minimum is achieved when the viewing angle is $0^{\circ}$ and the spin parameter approaches $1$. However, with a non-zero disk tilt, the inner shadow exhibits novel configurations–taking on petal, crescent, or eyebrow shapes–significantly smaller than $S_{\textrm{min}}$ across various parameter spaces. This indicates that the inner shadow is highly sensitive to the accretion environment, suggesting that caution is needed when using it as a diagnostic tool for black holes. Notably, an observed inner shadow smaller than $S_{\textrm{min}}$ would either indicate the presence of a tilted accretion disk or support the viability of modified gravity. Moreover, in certain parameter spaces, we identify the emergence of a dual-shadow structure, which could also serve as a probe for the tilted accretion disk.
@@ 1 Introduction
Since the advent of general relativity, black holes have remained a subject of profound fascination within the astronomical community. In recent years, the study of black holes has evolved from theoretical speculation to direct observation. Notably, the Event Horizon Telescope (EHT) collaboration, utilizing a Very Long Baseline Interferometry (VLBI) array, successfully captured and released the first images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way [1, 2]. These landmark observations not only conclusively verified the existence of black holes but also opened a novel channel for their exploration. Subsequently, a substantial body of research on black hole images [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], black hole-jet system images [18, 19, 20, 21, 22], hot-spot images [23, 24, 25, 26, 27, 28], and polarized images [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], within the frameworks of general relativity and modified gravity theories, has emerged, injecting unprecedented vitality into this field.
Theoretically, a black hole can capture light rays emitted by surrounding sources, resulting in a distinct region of brightness depression on the observation plane from Earth. This region is referred to as the ‘shadow’, first introduced by Falcke et al. [39], and its boundary is the 2D projection of the black hole’s unstable critical photon orbits. Notably, the black hole shadow is considered the ‘fingerprint’ of spacetime, unaffected by the accretion environment, making it a powerful tool for studying black hole physics and gravity theories. Consequently, many pioneering studies have focused on constraining system parameters through the shadow’s profile [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. However, accurately extracting the geometric information of the black hole shadow requires extremely high angular resolution, which is currently beyond reach with the technology of the EHT [4, 52]. As a result, despite the groundbreaking achievements of the EHT, there remains ample room for discussion regarding modifications to gravity theories and the accretion environment.
@@
In astrophysics, the angular momentum of accreting matter is not always aligned with the angular momentum of the black hole, which leads to the formation of tilted accretion disk. There is substantial evidence showing that tilted accretion disks are commonly present in X-ray binary systems [64, 65, 66], and the tilted accretion disk-jet system of the supermassive black hole M87 has also been confirmed [67]. Research on the structure evolution and energy spectrum of tilted accretion disks has garnered considerable attention in the scientific community [68, 69, 70, 71, 72]. Inspired by this, Hu and colleagues conducted a simulation of the images of spherically symmetric hairy black holes illuminated by a geometrically thin, tilted accretion disk within the framework of Horndeski theory [14]. They found that the inner shadow of the black hole tends to be obscured by the tilted accretion disk, and it also rotates along with the precession of the disk. Therefore, they argued that, given the difficulties in constraining the accretion environment, using the inner shadow to test gravity theories requires particular caution. However, the scientific community has reached a consensus on the existence of rotation in astrophysical black holes, indicating that Hu et al. did not tell the whole story. In the scenario of a rotating black hole, does a tilted accretion disk obscure the inner shadow? What insights can the inner shadow provide for black hole physics and high-energy physics? These questions will be properly addressed in this paper.
The remainder of this paper is organized as follows. In section 2, we briefly review the Kerr black hole and the canonical equation of photon motion, which form the foundation of the ray-tracing algorithm. Then, in section 3, we provide a detailed introduction to the backward ray-tracing method and the model of the tilted accretion disk. Before the conclusions and discussion in section 5, we present in section 4 the novel inner shadows cast by the Kerr black hole illuminated by a tilted thin accretion disk across different parameter spaces, and carefully elaborate on their astrophysical implications. Throughout this paper, we adopt geometric units, where the black hole mass $M$, the speed of light $c$, and the gravitational constant $G$ are set to unity.
@@ 3 Ray-tracing method
As shown in figure 1, the local frames of the observer and the Kerr black hole are represented by quasi-Cartesian coordinate systems $(x,y,z)$ and $(x^{\prime},y^{\prime},z^{\prime})$, respectively. Here, the $z$-axis coincides with $\overline{oo^{\prime}}$ and points towards the point $o^{\prime}$, with an angle $\Theta$ relative to the black hole’s rotation axis ($z^{\prime}$-axis), which defined as the viewing angle. The observer’s screen, depicted as a purple block, lies in $\overline{xoy}$ plane, thus $(x,y,0)$ naturally denotes the initial position of the photon. The solid disk inclined at an angle $\sigma$ to the black hole’s equatorial plane represents a tilted, geometrically thin accretion disk. It is divided into four regions of different colors by quadrants to vividly illustrate the black hole’s inner shadow and the spacetime distortion. It is important to note that for the equatorial accretion disk scenario $(\sigma=0^{\circ})$, the black hole image is independent of the observer’s azimuthal angle. However, this property is expected to not hold when a non-zero $\sigma$ is introduced. Therefore, an additional parameter, the observation azimuth $\Phi$, is included in our model.
@@
Now, we have obtained the initial conditions of the photon in the local frame of the black hole as $(t,r,\theta,\phi,p_{t},p_{r},p_{\theta},p_{\phi})$. Using a fifth- and sixth-order Runge-Kutta-Fehlberg integrator with adaptive step size (RKF56), we can trace the propagation of the photon corresponding to each pixel on the observation plane, thereby simulating the black hole’s inner shadow.
It is worth mentioning that, to improve the efficiency of the ray-tracing method, we supplemented the RKF56 with an orbit-adaptive step size adjustment, which significantly enhances the convergence speed of the algorithm. Specifically, the iteration step size $\tau$ in the algorithm is dynamically adjusted based on the distance between the photon and the black hole’s event horizon, $\tau=\tau_{0}(r/r_{\textrm{e}})^{\kappa}$. In this paper, we set $\tau_{0}$ and $\kappa$ to $10^{-5}$ and $1.8$, respectively.
@@ 4 Results
In this section, we fix the observation distance at $r_{\textrm{obs}}=1000$ M and the field of view at $16\times 16$ M with a resolution of $400\times 400$ pixels to numerically simulate the inner shadow of a Kerr black hole illuminated by a tilted thin accretion disk across different parameter spaces. It is worth noting that, in the case of an equatorial accretion disk, the black hole image remains independent of the observer’s azimuth $\Phi$. However, this invariance does not hold for an inclined accretion disk scenario. Therefore, we present the results for $\Phi=0^{\circ}$ and $\Phi=90^{\circ}$ separately and analyze the impact of the accretion environment on the inner shadow.
@@ 4.1 $\Phi=0^{\circ}$ case
When the observer’s azimuth is fixed at zero, we numerically simulate the inner shadows of the Kerr black holes illuminated by a tilted thin accretion disk with varying inclinations for observation angle of $17^{\circ}$, $50^{\circ}$, and $85^{\circ}$. The results are presented in figures 2-4, respectively.
Figure 2: Evolution of the inner shadow of a Kerr black hole illuminated by a tilted thin accretion disk with respect to disk inclination $\sigma$ and spin parameter $a$. From left to right, the values of disk inclination are $0^{\circ}$, $15^{\circ}$, $30^{\circ}$, $45^{\circ}$, and $60^{\circ}$; from top to bottom, the spin parameters are 00, $0.54$, $0.94$, and $0.9985$. Here, the observation angle and azimuth are fixed at $17^{\circ}$ and $0^{\circ}$, respectively, and the field of view is set to $16\times 16$ M, with a resolution of $400\times 400$ pixels. It is evident that the tilted accretion disk introduces a significant obscuration effect on the inner shadow.
@@
Additionally, the size of the inner shadow demonstrates a positive correlation with the observation angle and a negative correlation with the spin parameter. In other words, for a Kerr black hole endowed with an equatorial accretion disk, there exists a minimum area of the inner shadow, denoted as $S^{\textrm{min}}_{\Phi}$, for each observation angle. It is anticipated that as the spin parameter approaches $1$, $S^{\textrm{min}}_{\Phi=0^{\circ}}$ reaches its minimum value, corresponding to the smallest inner shadow area, $S_{\textrm{min}}$, that a Kerr black hole illuminated by an equatorial accretion disk can exhibit. We find that $S_{\textrm{min}}=13.075$ M2.
@@
Interestingly, we find that when the accretion disk deviates from the equatorial plane, the inner shadow of the black hole can shrink further from $S^{\textrm{min}}_{\Phi}$: the size of the inner shadow decreases with an increase in the tilt of the accretion disk, and this reduction is positively correlated with the black hole’s spin parameter, as vividly illustrated in the last row of figure 2. This suggests that if a significantly smaller than expected inner shadow is observed in reality, it may indicate the presence of a tilted accretion disk surrounding the black hole. In other words, the size of the inner shadow can serve as a probe for identifying tilted accretion disks.
Moreover, we note that the introduction of the accretion disk inclination can alter the shape of the black hole’s inner shadow. When the spin parameter is zero, changes in the inner shadow due to the disk inclination can be fully attributed to a specific observation angle, a degeneracy that was previously identified in our previous work [14]. However, for non-zero spin parameters, we do not observe a degeneracy between the observation angle and the disk tilt; instead, the presence of the tilted disk leads to a richer variety of inner shadow profiles, including crescent, eyebrow, and petal shapes. This phenomenon occurs because the tilted accretion disk obscures light rays that would otherwise cross the black hole’s event horizon, thereby affecting the inner shadow’s visibility.
@@
In figures 3(n) and (s), we identify escaping photons (depicted in gray) that, upon emission from the observation plane, neither get absorbed by the black hole nor collide with the accretion disk; instead, they deflect near the black hole and travel to infinity. Figure 5 illustrates the trajectory of the photon corresponding to the observation coordinate $(-3.46867,-1.82456)$ from figure 3(n) within the black hole’s local reference frame (a), along with its projections onto the $\overline{x^{\prime}o^{\prime}z^{\prime}}$ plane (b) and $\overline{x^{\prime}o^{\prime}y^{\prime}}$ plane (c). It is evident that the photon consistently remains below the tilted accretion disk throughout its trajectory, without making contact with the disk, as demonstrated in the zoomed-in view in (b). Notably, in the case of the equatorial plane, the occurrence of escaping photons is considerably less likely, as evidenced by the intersection of the photon trajectory with the green dashed-dot line in (b). For a tilted accretion disk of finite size, escaping photons can only manifest in specific parameter spaces, a phenomenon resulting from the combined effects of frame dragging and the tilted disk. What role does this type of light play in black hole observation that deserves further investigation.
@@
Figure 5: Trajectory of an escaping photon in the black hole’s local reference frame (a) and its projection onto the $\overline{x^{\prime}o^{\prime}z^{\prime}}$ plane (b) and the $\overline{x^{\prime}o^{\prime}y^{\prime}}$ plane (c). In panel (b), the tilted accretion disk ($\sigma=45^{\circ}$) and the equatorial plane are represented by the blue and green dashed-dot lines, respectively. It can be observed that the escaping photon neither crosses the black hole’s event horizon nor collides with the accretion disk. Instead, it escapes to infinity. Notably, this photon can intersect the equatorial accretion disk, as indicated by the intersection of the photon trajectory and the green dashed-dot line in (b). Here, the viewing angle is $50^{\circ}$, and the spin parameter is $0.94$.

### v3

# Novel inner shadows of the Kerr black hole with a tilted thin accretion disk

@@ Abstract
The inner shadow of a black hole, as a projection of the event horizon, is regarded as a potential tool for testing gravitational theories and constraining system parameters. Whether this holds in the case of a tilted accretion disk warrants further investigation. In this paper, we employ a ray-tracing algorithm to simulate images of the Kerr black hole with a tilted thin accretion disk, with particular attention to the relationship between the inner shadow and system parameters. Our findings reveal that in the case of an equatorial accretion disk, the Kerr black hole exhibits a minimum inner shadow size of $S_{\textrm{min}}=13.075$ M2, where M denotes the black hole mass. This minimum is achieved when the viewing angle is $0^{\circ}$ and the spin parameter approaches $1$. However, with a non-zero disk tilt, the inner shadow exhibits novel configurations–taking on petal, crescent, or eyebrow shapes–significantly smaller than $S_{\textrm{min}}$ across various parameter spaces. This indicates that the inner shadow is highly sensitive to the accretion environment, suggesting that caution is needed when using it as a diagnostic tool for black holes. Notably, an observed inner shadow smaller than $S_{\textrm{min}}$ would either indicate the presence of a tilted accretion disk or support the viability of modified gravity. Moreover, in certain parameter spaces, we identify the emergence of a dual-shadow structure, which could also serve as a probe for the tilted accretion disk.
@@ 1 Introduction
Since the advent of general relativity, black holes have remained a subject of profound fascination within the astronomical community. In recent years, the study of black holes has evolved from theoretical speculation to direct observation. Notably, the Event Horizon Telescope (EHT) collaboration, utilizing a Very Long Baseline Interferometry (VLBI) array, successfully captured and released the first images of the supermassive black holes at the centers of the M87 galaxy and the Milky Way [1, 2]. These landmark observations not only conclusively verified the existence of black holes but also opened a novel channel for their exploration. Subsequently, a substantial body of research on black hole shadows [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], black hole-jet system images [18, 19, 20, 21, 22], hot-spot images [23, 24, 25, 26, 27, 28], and polarized images [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], within the frameworks of general relativity and modified gravity theories, has emerged, injecting unprecedented vitality into this field.
Theoretically, a black hole can capture light rays emitted by surrounding sources, resulting in a distinct region of brightness depression on the observation plane from Earth. This region is referred to as the ‘shadow’, first introduced by Falcke et al. [39], and its boundary is the 2D projection of the black hole’s unstable critical photon orbits. Notably, the black hole shadow is considered the ‘fingerprint’ of spacetime, unaffected by the accretion environment, making it a powerful tool for studying black hole physics and gravity theories. Consequently, many pioneering studies have focused on constraining system parameters through the shadow’s profile [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. However, accurately extracting the geometric information of the black hole shadow requires extremely high angular resolution, which is currently beyond reach with the technology of the EHT [4, 55]. As a result, despite the groundbreaking achievements of the EHT, there remains ample room for discussion regarding modifications to gravity theories and the accretion environment.
@@
In astrophysics, the angular momentum of accreting matter is not always aligned with the angular momentum of the black hole, which leads to the formation of tilted accretion disk. There is substantial evidence showing that tilted accretion disks are commonly present in X-ray binary systems [67, 68, 69], and the tilted accretion disk-jet system of the supermassive black hole M87 has also been confirmed [70]. Research on the structure evolution and energy spectrum of tilted accretion disks has garnered considerable attention in the scientific community [71, 72, 73, 74, 75]. Inspired by this, Hu and colleagues conducted a simulation of the images of spherically symmetric hairy black holes illuminated by a geometrically thin, tilted accretion disk within the framework of Horndeski theory [14]. They found that the inner shadow of the black hole tends to be obscured by the tilted accretion disk [^2] The tilted accretion disk acts as a light source, making its obscuration of the black hole’s inner shadow seemingly paradoxical. However, this is physically reasonable. From the perspective of backward ray-tracing, the tilted disk can intercept some photons that would otherwise fall into the black hole (In the case of an equatorial accretion disk, these photons would contribute to the inner shadow), resulting in non-zero luminosities for the corresponding observational pixels. In other words, regions that should appear dark (the inner shadow) are illuminated by the tilted accretion disk, while the inner shadow itself becomes obscured., and it also rotates along with the precession of the disk. Therefore, they argued that, given the difficulties in constraining the accretion environment, using the inner shadow to test gravity theories requires particular caution. However, the scientific community has reached a consensus on the existence of rotation in astrophysical black holes, indicating that Hu et al. did not tell the whole story. In the scenario of a rotating black hole, does a tilted accretion disk obscure the inner shadow? What insights can the inner shadow provide for black hole physics and high-energy physics? To address these issues, this study defines a tilted plane from a purely geometric perspective to emulate an inclined accretion disk surrounding a Kerr black hole. Employing a ray-tracing algorithm, we compute trajectories for a large number of light rays, identifying those intersecting either the tilted accretion disk or the black hole’s event horizon. This approach reveals intriguing features of the inner shadow induced by the disk inclination in rotating black holes.
The remainder of this paper is organized as follows. In section 2, we briefly review the Kerr black hole and the canonical equation of photon motion, which form the foundation of the ray-tracing algorithm. Then, in section 3, we provide a detailed introduction to the backward ray-tracing method and the model of the tilted accretion disk. Before the conclusions and discussion in section 5, we present in section 4 the novel inner shadows cast by the Kerr black hole surrounded by a tilted thin accretion disk across different parameter spaces, and carefully elaborate on their astrophysical implications. Throughout this paper, we adopt geometric units, where the black hole mass $M$, the speed of light $c$, and the gravitational constant $G$ are set to unity.
@@ 3 Ray-tracing method
As shown in figure 1, the local frames of the observer and the Kerr black hole are represented by quasi-Cartesian coordinate systems $(x,y,z)$ and $(x^{\prime},y^{\prime},z^{\prime})$, respectively. Here, the $z$-axis coincides with $\overline{oo^{\prime}}$ and points towards the point $o^{\prime}$, with an angle $\Theta$ relative to the black hole’s rotation axis ($z^{\prime}$-axis), which defined as the viewing angle. The observer’s screen, depicted as a purple block, lies in $\overline{xoy}$ plane, thus $(x,y,0)$ naturally denotes the initial position of the photon. The solid disk inclined at an angle $\sigma$ to the black hole’s equatorial plane represents a tilted, geometrically thin accretion disk. It is divided into four regions of different colors by quadrants to vividly illustrate the black hole’s inner shadow and the spacetime distortion. Consequently, for each set of system parameters, the observation plane will exhibit regions colored red, yellow, blue, and green contributed by rays hitting the tilted plane, along with black regions corresponding to rays that fall into the black hole. It is important to note that for the equatorial accretion disk scenario $(\sigma=0^{\circ})$, the black hole image is independent of the observer’s azimuthal angle. However, this property is expected to not hold when a non-zero $\sigma$ is introduced. Therefore, an additional parameter, the observation azimuth $\Phi$, is included in our model.
@@
Now, we have obtained the initial conditions of the photon in the local frame of the black hole as $(t,r,\theta,\phi,p_{t},p_{r},p_{\theta},p_{\phi})$. Using a fifth- and sixth-order Runge-Kutta-Fehlberg integrator with adaptive step size (RKF56), we can trace the propagation of the photon corresponding to each pixel on the observation plane, thereby simulating the black hole’s inner shadow. It is crucial to emphasize that when determining whether a photon crosses the black hole’s event horizon, achieving a photon radial coordinate exactly satisfying $r=r_{\textrm{e}}$ is numerically improbable, as this constitutes a coordinate singularity. To address this, we implement a numerical criterion: once a photon’s position meets $r\leq r_{\textrm{e}}+10^{-4}$, the photon is deemed to have entered the black hole, and the simulation for that particular ray is terminated.
Additionally, to improve the efficiency of the ray-tracing method, we supplemented the RKF56 with an orbit-adaptive step size adjustment, which significantly enhances the convergence speed of the algorithm. Specifically, the iteration step size $\tau$ in the algorithm is dynamically adjusted based on the distance between the photon and the black hole’s event horizon, $\tau=\tau_{0}(r/r_{\textrm{e}})^{\kappa}$. In this paper, we set $\tau_{0}$ and $\kappa$ to $10^{-5}$ and $1.8$, respectively.
@@ 4 Results
In this section, we fix the observation distance at $r_{\textrm{obs}}=1000$ M and the field of view at $16\times 16$ M with a resolution of $400\times 400$ pixels to numerically simulate the inner shadow of a Kerr black hole with a tilted thin accretion disk across different parameter spaces. It is worth noting that, in the case of an equatorial accretion disk, the black hole image remains independent of the observer’s azimuth $\Phi$. However, this invariance does not hold for an inclined accretion disk scenario. Therefore, we present the results for $\Phi=0^{\circ}$ and $\Phi=90^{\circ}$ separately and analyze the impact of the accretion environment on the inner shadow.
@@ 4.1 $\Phi=0^{\circ}$ case
When the observer’s azimuth is fixed at zero, we numerically simulate the inner shadows of the Kerr black holes with a tilted thin accretion disk with varying inclinations for observation angle of $17^{\circ}$, $50^{\circ}$, and $85^{\circ}$. The results are presented in figures 2-4, respectively.
Figure 2: Evolution of the inner shadow of a Kerr black hole with a tilted thin accretion disk with respect to disk inclination $\sigma$ and spin parameter $a$. From left to right, the values of disk inclination are $0^{\circ}$, $15^{\circ}$, $30^{\circ}$, $45^{\circ}$, and $60^{\circ}$; from top to bottom, the spin parameters are 00, $0.54$, $0.94$, and $0.9985$. Here, the observation angle and azimuth are fixed at $17^{\circ}$ and $0^{\circ}$, respectively, and the field of view is set to $16\times 16$ M, with a resolution of $400\times 400$ pixels. It is evident that the tilted accretion disk introduces a significant obscuration effect on the inner shadow.
@@
Additionally, the size of the inner shadow demonstrates a positive correlation with the observation angle and a negative correlation with the spin parameter. In other words, for a Kerr black hole endowed with an equatorial accretion disk, there exists a minimum area of the inner shadow, denoted as $S^{\textrm{min}}_{\Theta}$, for each observation angle. It is anticipated that as the spin parameter approaches $1$, $S^{\textrm{min}}_{\Theta=0^{\circ}}$ reaches its minimum value, corresponding to the smallest inner shadow area, $S_{\textrm{min}}$, that a Kerr black hole surrounded by an equatorial accretion disk can exhibit. We find that $S_{\textrm{min}}=13.075$ M2.
@@
Interestingly, we find that when the accretion disk deviates from the equatorial plane, the inner shadow of the black hole can shrink further from $S^{\textrm{min}}_{\Theta}$: the size of the inner shadow decreases with an increase in the tilt of the accretion disk, and this reduction is positively correlated with the black hole’s spin parameter, as vividly illustrated in the last row of figure 2. This suggests that if a significantly smaller than expected inner shadow is observed, it may indicate the presence of a tilted accretion disk surrounding the black hole. In other words, the size of the inner shadow can serve as a probe for identifying tilted accretion disks.
Moreover, we note that the introduction of the accretion disk inclination can alter the shape of the black hole’s inner shadow. When the spin parameter is zero, changes in the inner shadow due to the disk inclination can be fully attributed to a specific observation angle, a degeneracy that was identified in our previous work [14]. However, for non-zero spin parameters, we do not observe a degeneracy between the observation angle and the disk tilt; instead, the presence of the tilted disk leads to a richer variety of inner shadow profiles, including crescent, eyebrow, and petal shapes. This phenomenon, along with the inner shadow shrinkage reported earlier, occurs because the tilted accretion disk blocks light rays that would otherwise cross the event horizon, thus reducing the inner shadow’s visibility. To better visualize this characteristic, we employ ray propagation to demonstrate how the tilted accretion disk erodes the black hole’s inner shadow. As shown in figure 5(a), a Kerr black hole with spin parameter $0.94$ is represented by a black sphere, while the accretion disk with inclination $\sigma=30^{\circ}$ is depicted as a four-colored tilted plane, and the equatorial disk appears as a gray disk. The purple, blue, and yellow curves correspond to three light rays originating from observational coordinates $(2,1)$, $(-2,0)$, and $(1,0.2)$, respectively. Clearly, in the equatorial disk scenario, all three light rays can plunge into the black hole without intersecting the accretion disk. However, when the disk is tilted to $30^{\circ}$, it intercepts the blue and purple rays before they reach the black hole. From an astrophysical standpoint, this implies that while these rays would contribute to the dark regions of the observation plane for an equatorial disk, in the tilted configuration they instead carry disk radiation, resulting in non-zero luminosities at the observational coordinates $(2,1)$ and $(-2,0)$. By contrast, the yellow ray consistently contributes to the black hole’s inner shadow, as it directly intersects the event horizon in both disk configurations. Figures 5(b) and (c) display the projections of the three light rays from panel (a) in the $\overline{x^{\prime}o^{\prime}y^{\prime}}$ and $\overline{x^{\prime}o^{\prime}z^{\prime}}$ planes, respectively. A clear distinction emerges: when the accretion disk lies in the equatorial plane (green dashed line in the right panel), all three rays bypass the disk and plunge directly into the black hole, thereby contributing to the inner shadow. However, when the disk rotates to the red dashed position, it intercepts both the blue and purple rays, inducing deformation of the inner shadow.
@@
Figure 5: Trajectories of three photons in the black hole’s local reference frame (a) and their projections onto the (b) $\overline{x^{\prime}o^{\prime}y^{\prime}}$ and $\overline{x^{\prime}o^{\prime}z^{\prime}}$ planes. In panel (a), the equatorial accretion disk is represented by a gray disk, while the black hole’s event horizon is shown as a black sphere. Panels (b) and (c) display the event horizon as a gray disk, with the $30^{\circ}$-tilted accretion disk and equatorial disk marked by red and green dashed lines, respectively. The results demonstrate that under the equatorial disk configuration, all three rays (blue, purple, and yellow) contribute to the inner shadow, as they reach the event horizon without intersecting the disk. As evident in panel (c), these photons maintain trajectories entirely above the equatorial plane. However, when the disk is tilted to $30^{\circ}$, the blue and purple rays intersect the inclined plane before entering the black hole, indicating that their corresponding regions on the observational screen will exhibit non-zero luminosity. This reveals that the tilted accretion disk effectively illuminates portions of the black hole’s inner shadow. Here, we have the viewing angle of $17^{\circ}$, and the spin parameter of $0.94$.
@@
In figures 3(n) and (s), we identify escaping photons (depicted in gray) that, upon emission from the observation plane, neither get absorbed by the black hole nor collide with the accretion disk; instead, they deflect near the black hole and travel to infinity. Figure 6 illustrates the trajectory of the photon corresponding to the observational coordinate $(-3.46867,-1.82456)$ from figure 3(n) within the black hole’s local reference frame (a), along with its projections onto the $\overline{x^{\prime}o^{\prime}z^{\prime}}$ plane (b) and $\overline{x^{\prime}o^{\prime}y^{\prime}}$ plane (c). It is evident that the photon consistently remains below the tilted accretion disk throughout its trajectory, without making contact with the disk, as demonstrated in the zoomed-in view in (b). Notably, in the case of the equatorial plane, the occurrence of escaping photons is considerably less likely, as evidenced by the intersection of the photon trajectory with the green dashed line in (b). For a tilted accretion disk of finite size, escaping photons can only manifest in specific parameter spaces, a phenomenon resulting from the combined effects of frame dragging and the tilted disk.
@@
Figure 6: Trajectory of an escaping photon in the black hole’s local reference frame (a) and its projection onto the $\overline{x^{\prime}o^{\prime}z^{\prime}}$ plane (b) and the $\overline{x^{\prime}o^{\prime}y^{\prime}}$ plane (c). In panel (b), the tilted accretion disk ($\sigma=45^{\circ}$) and the equatorial plane are represented by the blue and green dashed lines, respectively. It can be observed that the escaping photon neither crosses the black hole’s event horizon nor collides with the accretion disk. Instead, it escapes to infinity. Notably, this photon can intersect the equatorial accretion disk, as indicated by the intersection of the photon trajectory and the green dashed line in (b). Here, the viewing angle is $50^{\circ}$, and the spin parameter is $0.94$.
It should be emphasized that the inner shadow under tilted disk conditions studied in this work constitutes a rigorous generalization of the black hole’s inner shadow in the equatorial disk scenario [6]. Specifically, the inner shadow boundaries in figures 2-4 represent the direct image of the intersection between the event horizon and the tilted plane of the accretion disk. In other words, the inner shadow studied in this work is defined purely geometrically: the boundary corresponds to the projection onto the observation screen of the intersection curve between an arbitrarily inclined plane and the black hole’s event horizon. The formation of such an inner shadow requires two conditions to be met. First, the accretion disk’s inner boundary must extend to the black hole’s event horizon. Second, from the perspective of backward ray tracing, photons contributing to the inner shadow must intersect the event horizon without crossing the accretion disk. Figures 7 and 8 display the number of intersections between light rays and the tilted accretion disk for each pixel on the observation plane across a broad parameter space. Here, the black color indicates zero intersections, meaning the light rays bypass the accretion disk and directly cross the event horizon, thereby contributing to the inner shadow. Indeed, one can clearly observe that the black regions in figures 7 and 8 exhibit coincidence with the inner shadows shown in the corresponding panels of figures 2 and 4. In the equatorial accretion disk scenario (left columns of figures 7 and 8), rays intersecting the disk multiple times contribute to the lensing band (yellow regions). Notably, these lensing bands progressively approach the critical curve [4] as the intersection count increases, as demonstrated by the green curves within the bands. By contrast, such higher-order image convergence behavior is rarely observed in the tilted disk configuration. We find that for smaller black hole spins and disk inclination angles (i.e., panels (g) and (h) in figure 7), multiply disk-crossing rays primarily concentrate near the critical curve, though their signatures remain identifiable near the inner shadow. However, as both $\sigma$ and $a$ increase, these higher-order images become distributed across multiple distinct regions without convergence behavior, as evidenced by the green, blue, and gray patterns in panels (l)-(o) and (q)-(t) of figure 7. This phenomenon occurs because a tilted accretion disk effectively introduces additional photon-intercepting planes on both sides of the equatorial plane, substantially increasing the probability of ray-disk interactions. This effect is particularly pronounced for photons orbiting near the equatorial plane under frame-dragging effects, which can rapidly accumulate multiple disk intersections. However, such intersection multiplicity does not provide meaningful guidance for identifying the black hole’s characteristic geometry or lensing bands. Nevertheless, from an astrophysical perspective, the multiple intersections between light rays and the tilted accretion disk can enhance the specific intensities of these photons, producing bright features in the image. This phenomenon warrants further investigation in future studies.
Figure 7: Distribution of light ray-disk intersection counts on the observation plane. From left to right, the values of disk inclination are $0^{\circ}$, $15^{\circ}$, $30^{\circ}$, $45^{\circ}$, and $60^{\circ}$; from top to bottom, the spin parameters are 00, $0.54$, $0.94$, and $0.9985$. Here, the observation angle and azimuth are fixed at $17^{\circ}$ and $0^{\circ}$, respectively. Light rays intersecting the accretion disk 00, $1$, $2$, $3$, $4$, and $5$ times are represented by black, red, yellow, green, blue, and gray, respectively. Notably, in the tilted disk scenario, photons contributing to the black hole’s inner shadow must have zero intersections with the disk. Moreover, compared to the equatorial case (left column), images formed by multiply disk-crossing rays under tilted disk conditions generally lack asymptotic convergence.